# DPNet: Doppler LiDAR Motion Planning for Highly-Dynamic Environments

### v1

@@ Abstract
Existing motion planning methods often struggle with rapid-motion obstacles due to an insufficient understanding of environmental changes. To address this limitation, we propose integrating motion planners with Doppler LiDARs which provide not only ranging measurements but also instantaneous point velocities. However, this integration is nontrivial due to the dual requirements of high accuracy and high frequency. To this end, we introduce Doppler Planning Network (DPNet), which tracks and reacts to rapid obstacles using Doppler model-based learning. Particularly, we first propose a Doppler Kalman neural network (D-KalmanNet) to track the future states of obstacles under partially observable Gaussian state space model. We then leverage the estimated motions to construct a Doppler-tuned model predictive control (DT-MPC) framework for ego-motion planning, enabling runtime auto-tuning of the controller parameters. These two model-based learners allow DPNet to maintain lightweight while learning fast environmental changes using minimum data, and achieve both high frequency and high accuracy in tracking and planning. Experiments on both high-fidelity simulator and real-world datasets demonstrate the superiority of DPNet over extensive benchmark schemes. Project page: https://github.com/UUwei-zuo/DPNet.git
@@
Model-based learning, motion planning
@@ I Introduction
Real-time robot motion planning in highly-dynamic environments is crucial for a broad range of applications, including emergency rescue and autonomous driving [han2025hierarchically]. Such scenarios require accurate knowledge of environmental dynamics, which can be leveraged by the motion controller to generate sequences of collision-free actions under kinematic constraints. Existing motion planning methods heavily rely on light detection and ranging (LiDAR) sensors to capture the current positions of the obstacles from the point cloud [flores2018cooperative], and then predict obstacles’ future trajectories via learning-based [milan2017online], physics-based [lefkopoulos2020interaction], or physics-informed learning [revach2022kalmannet, liao2024physics] methods. But these approaches struggle in handling rapid-motion obstacles due to their limited ability to understand fast environmental changes in real time.
@@
Recently, however, the situation has changed thanks to the remarkable milestone of the emerging Doppler LiDAR sensors, which not only offer the high-resolution point range data but also the point Doppler velocity [hexsel2022dicp]. The new dimension of Doppler velocity explicitly provides the instantaneous motion information of obstacles, opening up new opportunities for agile motion planning in rapid-changing scenarios. Yet, how to integrate Doppler LiDAR sensors into motion planners for a unified solution framework is still unknown. This is nontrivial since the solution should be high-performance, real-time, safety-assured, and work properly with limited onboard computational resources.
@@
Figure 1: Doppler model-based learning enables DPNet to understand and traverse highly-dynamic environments.
@@
Given a certain estimate of the obstacle state, the next step is to conduct motion control. Conventional motion controllers, e.g., model predictive control (MPC), use the estimated dynamics to construct consecutive problems along the horizon [zhang2024online]. However, they might not work if the prior knowledge about the cost function or the dynamical model is imperfect. To address this problem, emerging auto-tuning motion controllers enable parameter learning from closed-loop execution data [zuliani2024closed]. Nonetheless, when applied to dynamic collision avoidance, these tune-from-testing methods typically require an actual collision event to trigger parameter update. To broaden their applications, how to realize run-time controller tuning without relying on actual collisions becomes an important issue.
To fill these research gaps, this paper proposes Doppler planning network (DPNet) shown in Fig. 1, which leverages the new dimension provided by Doppler LiDAR and two key algorithm innovations to achieve environmental motion learning and runtime controller tuning. The first innovation is a Doppler Kalman neural network (D-KalmanNet) which achieves real-time obstacle future state tracking under partially observable Gaussian state space (GSS) model. The D-KalmanNet incorporates the structural GSS model with a recurrent neural network (RNN) and sequences of Doppler velocity measurements, ensuring very high tracking frequency with an ultra-lightweight model size. The second innovation is a Doppler-tuned model predictive control (DT-MPC) framework design, which first performs Doppler-inferred collision check along the future horizon and then tunes the controller parameters using the imagined collision cost. Under the Doppler guidance, the controller learning can be performed in an online fashion before actual collisions. To reduce the computational cost, a factorization-based method is designed to update controller parameters efficiently and an alternating direction method of multipliers (ADMM) algorithm is used to execute the controller. Built upon the above two innovations, DPNet is able to understand and cope with fast environmental changes using minimum data.
To evaluate DPNet with D-KalmanNet and DT-MPC algorithms, extensive robot operation system (ROS) experiments are performed. We show that DPNet outperforms state-of-the-art planners including RDA [han2023rda], MPC-D-CBF [jian2023dynamic], and OBCA [zhang2020optimization] on diverse metrics in Carla simulator [dosovitskiy2017carla]. Ablation studies confirm DT-MPC’s superiority over existing auto-tuning MPC, especially in highly-cluttered environments. Moreover, evaluations on real-world Doppler LiDAR dataset Aevascene [aevascenes] demonstrate that D-KalmanNet achieves lowest tracking errors over existing tracking methods including Kalman filter [kalman1960new] and KalmanNet [revach2022kalmannet] across wide execution frequency range (i.e., 1101–10 Hz), different horizon lengths (1101-10), and various road scenarios (i.e., Highway and City). It is also found that D-KalmanNet can be executed at 1510015-100 Hz on a low-cost Orin NX 16 GB chip when tracking up to 10 objects simultaneously. To the best of our knowledge, this is the first attempt to integrate Doppler LiDAR with motion planning. Our contributions are summarized as follows:
  • We introduce the Doppler LiDAR to support motion planning designs, resulting in a novel DPNet framework. By incorporating the new dimension of Doppler velocity for model-based learning, DPNet can proactively avoid highly-dynamic obstacles.
  • We propose D-KalmanNet for obstacle tracking, achieving higher accuracy than existing physical, learning, or hybrid methods, while simultaneously maintaining ultra-low computation cost and high frequency.
  • We propose DT-MPC to achieve run-time controller tuning by leveraging Doppler-inferred collision instead of learning from actual collisions.
  • We implement the complete DPNet framework, including both D-KalmanNet and DT-MPC algorithms. Extensive results show significant improvements over existing schemes in terms of diverse metrics.
II Related Work
@@
Doppler LiDAR. In contrast to traditional LiDAR [flores2018cooperative], Doppler LiDAR reshapes the landscape of robotic solutions by providing Doppler velocity measurements. It has been used in object detection, state estimation, and simultaneous localization and mapping [shi2025pod, gu2022learning, peng2021detection, wu2022picking, zhao2024fmcw, yoon2023need, zhao2024free]. In terms of evaluation, Carla [dosovitskiy2017carla] provides realistic Doppler LiDAR simulations [carla-aeva, hexsel2022dicp], and AevaScenes [aevascenes] dataset provides on-road Doppler LiDAR recordings. However, all existing research focuses on open-loop perception usages, and integrating Doppler LiDAR into closed-loop motion planning remains unexplored.
@@
Obstacle Tracking. Incorporating forecasts of obstacle future trajectories can significantly enhance planning performance. Currently, obstacle tracking can be divided into learning-based, physics-based, and hybrid methods. Learning-based methods require large datasets and substantial computational resources for training [zhou2025robust, zhang2022learning]. Physics-based methods (e.g., Kalman filter [lefkopoulos2020interaction]) capitalize on prior knowledge for high computational efficiency, while relying on predefined kinematic models that introduce inevitable bias and limit the ability to handle complex motions, especially when the state transition model is partially observed. Lastly, hybrid methods combine strengths of both categories, resulting in efficient and adaptable physics-informed learning solutions. For example, the recent KalmanNet [revach2022kalmannet] mitigates the drawbacks of Kalman filters by learning the Kalman gain from real data, so as to maintain both high inference accuracy and lightweight model size. This work extends physics-informed learning to Doppler LiDAR, which has not been studied yet.
@@
Motion Planning. Many modern motion planning systems (e.g., robot, autonomous vehicle) are based on MPC, which generates optimal control outputs in response to environmental inputs within a receding horizon window [shi2021advanced, han2025neupan]. By incorporating obstacle tracking, MPC can construct time-varying constraints and handle dynamic environments [zhang2024online]. In general, the MPC performance depends on the problem modeling. To find the best setting, auto-tuning MPC [cheng2024difftune, tao2024difftune] uses closed-loop feedback to optimize controller parameters. However, it typically involves time-consuming data collections and requires actual collision events to trigger parameter learning. In contrast, our proposed DT-MPC achieves run-time MPC tuning by leveraging Doppler-inferred collisions, thereby boosting the robustness for handling rapid-motion obstacles.
@@ III Problem Formulation
where $C_{t}(\mathcal{S}_{t})$ is the utility function, $d_{\text{safe}}$ is the safety distance, and $(\mathbb{G}_{h+1},{\mathbb{O}}_{h+1}^{n})$ are bounding boxes of ego-robot and the $n$-th obstacle, respectively. The minimum distance function between two bounding boxes is defined as:
$\mathbf{dist}(\mathbb{P},\mathbb{Q})=\min_{\mathbf{p},\mathbf{q}}\|\mathbf{p}-\mathbf{q}\|_{2},\ \text{s.t.}~\mathbf{p}\in\mathbb{P},\mathbf{q}\in\mathbb{Q}.$ (2)
Considering a reference path tracking task with target waypoints $\{\mathbf{s}_{h+1}^{\star}|h\in\mathcal{H}_{t}\}$, we formulate $C_{t}(\mathcal{S}_{t})=\sum_{h\in\mathcal{H}_{t}}\left\|\mathbf{s}_{h+1}-\mathbf{s}_{h+1}^{\star}\right\|_{2}^{2}$ [zhang2020optimization].
Problem $\mathsf{P}_{t}$ is nontrivial since estimating ${\mathbb{O}}_{h+1}^{n}$ results in inevitably biased $\hat{\mathbb{O}}_{h+1|t}^{n}$, which can significantly mismatch the actual ${\mathbb{O}}_{h+1}^{n}$ if obstacle $n$ has rapid motions. This error would further propagate to motion planning since $\mathbf{dist}(\mathbb{G}_{h+1},\hat{\mathbb{O}}_{h+1|t}^{n})\neq\mathbf{dist}(\mathbb{G}_{h+1},{\mathbb{O}}_{h+1}^{n})$.
@@ IV-A Doppler KalmanNet for Tracking
The ego robot uses a Doppler LiDAR to scan the surrounding environment and obtains $\mathcal{P}_{t}=\{\mathbf{c}_{t}^{i},\|\bm{\mu}_{t}^{i}\|_{2}\}_{i=1}^{I}$, where $I$ is the number of points in the frame, $\mathbf{c}^{i}_{t}\in\mathbb{R}^{3}$ and $\|\bm{\mu}^{i}_{t}\|_{2}\in\mathbb{R}$ are the spatial coordinate and the scalar Doppler velocity of the $i$-th point, respectively. Accordingly, we define the Doppler-augmented state $\mathbf{x}_{t}^{n}$ as
@@
where $(x_{t}^{n},y_{t}^{n})$, $\theta_{t}^{n}$, $\bm{v}_{t}^{n}$ and $\bm{a}_{t}^{n}$ represent the position, orientation, linear velocity and acceleration at the center, respectively. Each $\mathbf{x}_{t}^{n}$ can be used to generate a ${\mathbb{O}}_{t}^{n}$. Hence, we only need to estimate a sequence of $\{\hat{\mathbf{x}}_{t+1|t}^{n},\hat{\mathbf{x}}_{t+2|t}^{n},\ldots,\hat{\mathbf{x}}_{t+H|t}^{n}\}$ from $\mathcal{P}_{t}$. We can fuse $\mathcal{P}_{t}$ with a state transition model to achieve this goal, but two challenges exist. First, according to Doppler LiDAR manufacturers [aevascenes], real-world velocity measurements typically involve noises with standard deviation of $0.1$ m/s. Second, built upon past trajectories, the transition model may mismatch the actual obstacle motions in the future [revach2022kalmannet]. Our D-KalmanNet can simultaneously tackle both challenges.
@@
4:   Group Doppler LiDAR point cloud as $\mathcal{P}_{t}^{n}$
5:   Obtain $\mathbf{y}_{t}^{n}$ from Doppler Velocity Recification
6:   if $t=0$ then
7:    $\mathbf{x}_{0}^{n}\leftarrow\mathbf{y}_{0}^{n}$, $\hat{\mathbf{x}}_{1|0}^{n}\leftarrow\mathbf{T}\cdot\mathbf{x}_{0}^{n}$, $\hat{\mathbf{y}}_{1|0}^{n}\leftarrow\mathbf{U}\cdot\hat{\mathbf{x}}_{1|0}^{n}$
8:   else
9:    $\mathcal{K}_{t}^{n}\leftarrow\textsf{RNN}(\mathbf{x}^{n}_{t-1},\hat{\mathbf{x}}_{t|t-1}^{n},\hat{\mathbf{y}}_{t|t-1}^{n},\mathbf{y}_{t}^{n})$
10:    $\mathbf{x}_{t}^{n}\leftarrow\hat{\mathbf{x}}_{t|t-1}^{n}+\mathcal{K}_{t}^{n}\cdot(\mathbf{y}_{t}^{n}-\hat{\mathbf{y}}_{t|t-1}^{n})$
11:    $\hat{\mathbf{x}}_{h+1|t}^{n}\leftarrow\mathbf{T}^{h-t+1}\cdot\mathbf{x}_{t}^{n}$, $\forall h\in\mathcal{H}_{t}$
12:    $\hat{\mathbf{y}}_{t+1|t}^{n}\leftarrow\mathbf{U}\cdot\hat{\mathbf{x}}_{t+1|t}^{n}$
13:   end if
@@
15:  Generate $\hat{\mathcal{O}}_{t}^{n}$ from $\{\hat{\mathbf{x}}_{t+1|t}^{n},\hat{\mathbf{x}}_{t+2|t}^{n},\ldots,\hat{\mathbf{x}}_{t+H|t}^{n}\}$
16:  return $\hat{\mathcal{O}}_{t}^{n}$ and $t\leftarrow t+1$
17:end while
Figure 3: Doppler velocity rectification.
IV-A1 Doppler Velocity Rectification
$\bm{\breve{\mu}}_{t}^{i}=\|\bm{\mu}_{t}^{i}\|_{2}\cos\rho_{t}^{i}\frac{(\mathbf{I}-\mathbf{e}\mathbf{e}^{\top})(\mathbf{c}^{\text{D}}_{t}-\mathbf{c}_{t}^{i})}{\|(\mathbf{I}-\mathbf{e}\mathbf{e}^{\top})(\mathbf{c}^{\text{D}}_{t}-\mathbf{c}_{t}^{i})\|_{2}},$ (4)
With the above vector projection, we conduct Doppler LiDAR measurement grouping which clusters points within the obstacle’s bounding box as
$\mathcal{P}_{t}^{n}=\{\{\mathbf{c}_{t}^{i},\|\bm{\mu}_{t}^{i}\|_{2}\}|\mathbf{c}_{t}^{i}\sqsubseteq\mathbb{O}^{n}_{t}\}=\{p_{t}^{i_{n}}\}_{i_{n}=1}^{I_{n}},$ (5)
@@ IV-A1 Doppler Velocity Rectification
Algorithm 2 Doppler Collision Check and Tuning
1: Initialize $\mathcal{S}_{t-1}=\emptyset$
2:while True do
3:  $\varrho_{n}\leftarrow\infty$ for $n\in\mathcal{N}$
4:  $\kappa\leftarrow\kappa^{\text{init}}$
5:  Obtain $\hat{\mathcal{O}}_{t}^{n}$ from the D-KalmanNet
6:  for $h\in\mathcal{H}_{t-1}^{-}$ do
7:   for $n\in\mathcal{N}$ do
8:    if $\varrho_{n}=\infty$ and $\mathbf{dist}(\mathbb{G}^{\circ}_{h+1},\hat{\mathbb{O}}_{h+1|t}^{n})\leq d_{0}$ then
9:     $\varrho_{n}\leftarrow\kappa$
10:    end if
11:   end for
12:   Update $\kappa\leftarrow\kappa+\Delta\kappa$
13:  end for
14:  $t\leftarrow t+1$
15:end while

### v2

@@ Abstract
Existing motion planning methods often struggle with rapid-motion obstacles due to an insufficient understanding of environmental changes. To address this, we propose integrating motion planners with Doppler LiDARs, which provide not only ranging measurements but also instantaneous point velocities. However, this integration is nontrivial due to the requirements of high accuracy and high frequency. To this end, we introduce Doppler Planning Network (DPNet), which tracks and reacts to rapid obstacles via Doppler model-based learning. We first propose a Doppler Kalman neural network (D-KalmanNet) to track obstacle states under a partially observable Gaussian state space model. We then leverage the predicted motions of obstacles to construct a Doppler-tuned model predictive control (DT-MPC) framework for ego-motion planning, enabling runtime auto-tuning of controller parameters. These two modules allow DPNet to learn fast environmental changes from minimal data while remaining lightweight, achieving high frequency and high accuracy in both tracking and planning. Experiments on high-fidelity simulator and real-world datasets demonstrate the superiority of DPNet over extensive benchmark schemes.
@@
Integrated planning and learning, collision avoidance, Doppler motion planning
@@ I Introduction
Real-time robot motion planning in highly-dynamic environments is crucial for a broad range of applications, including emergency rescue and autonomous driving [8]. Such scenarios require accurate knowledge of environmental dynamics, which can be leveraged by the motion controller to generate sequences of collision-free actions under kinematic constraints. Existing motion planning methods heavily rely on light detection and ranging (LiDAR) sensors to capture the current positions of the obstacles from the point cloud [4], and predict obstacles future trajectories via learning-based [13], physics-based [11], or physics-informed learning [17, 12] methods. Nonetheless, these approaches struggle to handle rapid-motion obstacles due to their limited ability to understand real-time environmental changes, such as instantaneous velocities.
@@
Recently, however, the situation has changed thanks to the remarkable milestone of the emerging Doppler LiDAR sensors, which not only offer traditional point ranging data but also point Doppler velocity [9]. The additional dimension explicitly provides instantaneous obstacle motion information, opening up new opportunities for motion planning in rapidly changing scenarios. Yet, how to integrate Doppler LiDAR into motion planners within a unified solution framework remains an open challenge. This is nontrivial since the solution should be high-performance, real-time, safety-assured, and work properly with limited onboard computational resources.
@@
Figure 1: Comparison between DPNet and RDA [7]. DPNet can understand and traverse highly-dynamic environments safely.
Given certain predictions of obstacle future states, the next step is motion control. Conventional motion controllers, e.g., model predictive control (MPC), use the predicted states to construct consecutive problems along the horizon [28]. However, they may suffer from suboptimality if the prior knowledge about the cost function or the dynamics model is imperfect. To address this, emerging auto-tuning MPC enables parameter learning from closed-loop execution data [22]. Nevertheless, when applied to collision avoidance, these tune-from-execution methods typically require actual collision events to trigger the process. To broaden their applications, how to leverage Doppler information to trigger runtime tuning without relying on actual events becomes an important issue.
To fill these research gaps, this paper proposes Doppler planning network (DPNet). As shown in Fig. 1, DPNet can traverse highly-dynamic areas by leveraging Doppler LiDAR and introducing two key algorithmic innovations for obstacle motion learning and runtime controller tuning. The first innovation is a Doppler Kalman neural network (D-KalmanNet) which achieves real-time obstacle future state prediction under a partially observable Gaussian state space (GSS) model. The D-KalmanNet incorporates the structural GSS model with a recurrent neural network (RNN) and sequences of Doppler velocity measurements, ensuring a very high tracking frequency and a lightweight model size. The second innovation is a Doppler-tuned model predictive control (DT-MPC) framework that performs Doppler-inferred collision check to trigger runtime parameter tuning from imagined collisions instead of actual ones. For computational efficiency, a heuristic update strategy is designed and an alternating direction method of multipliers (ADMM) algorithm is used for controller execution. Built upon these innovations, DPNet can understand and cope with rapid environmental changes using minimal data.
To evaluate DPNet with D-KalmanNet and DT-MPC algorithms, extensive robot operation system (ROS) experiments are performed. We show that DPNet outperforms state-of-the-art planners including RDA [7], MPC-D-CBF [10], and OBCA [27] on diverse metrics in Carla simulator [3]. Ablation study and sensitivity analysis confirm DT-MPC’s effectiveness and robustness in triggering parameter updates without actual events, especially in cluttered environments. Moreover, evaluations on real-world Doppler LiDAR dataset Aevascene [15] demonstrate that D-KalmanNet achieves lowest tracking errors over existing tracking methods including Kalman filter [10] and KalmanNet [17] across wide execution frequency range (i.e., 1101–10 Hz), different horizon lengths (1101-10), and various road scenarios (i.e., Highway and City). It is also found that D-KalmanNet can be executed at 1510015-100 Hz on a low-cost NVIDIA Jetson Orin NX 16 GB when tracking up to 10 objects simultaneously. To the best of our knowledge, this is the first attempt to integrate Doppler LiDAR with motion planning. Our contributions are summarized as follows:
  • We introduce Doppler LiDAR to support motion planner designs, resulting in a novel DPNet framework. By incorporating Doppler velocities for obstacle prediction, DPNet can proactively avoid highly-dynamic obstacles.
  • We propose D-KalmanNet for obstacle state tracking and prediction by model-based learning, achieving higher accuracy than existing methods while maintaining low computational cost and high frequency.
  • We propose DT-MPC for runtime controller tuning by triggering updates from Doppler-inferred collisions.
  • We implement the complete DPNet framework for evaluation. Extensive results show significant improvements over existing schemes in terms of diverse metrics.
II Related Work
@@
Doppler LiDAR. In contrast to traditional LiDAR [4], Doppler LiDAR reshapes the landscape of robotic solutions by providing Doppler velocity measurements. It has been used in object detection, state estimation, and simultaneous localization and mapping [20, 5, 16, 23, 29, 24, 30] In terms of evaluation, Carla [3] provides realistic Doppler LiDAR simulations [1, 9], and AevaScenes [15] dataset provides on-road Doppler LiDAR recordings. However, all existing research focuses on open-loop perception usages, and integrating Doppler LiDAR into closed-loop motion planning remains unexplored.
@@
Obstacle Tracking. Incorporating forecasts of obstacle future trajectories can significantly enhance planning performance. Currently, obstacle tracking can be divided into learning-based, physics-based, and hybrid methods. Learning-based methods require large datasets and substantial computational resources for training [31, 25]. Physics-based methods (e.g., Kalman filter [11]) capitalize on prior knowledge for high computational efficiency, while relying on predefined kinematic models that introduce inevitable bias and limit the ability to handle complex motions, especially when the state transition model is partially observed. Lastly, hybrid methods combine strengths of both categories, resulting in efficient and adaptable physics-informed learning solutions. For example, the recent KalmanNet [17] mitigates the drawbacks of Kalman filters by learning the Kalman gain from real data, so as to maintain both high inference accuracy and lightweight model size. This work extends physics-informed learning to Doppler LiDAR, which has not been studied yet.
@@
Motion Planning. Many modern motion planning systems (e.g., robot, autonomous vehicle) are based on MPC, which generates optimal control outputs in response to environmental inputs within a receding horizon window [19]. By incorporating obstacle tracking, MPC can construct time-varying constraints and handle dynamic environments [28]. In general, the MPC performance depends on the problem modeling. To find the best setting, emerging auto-tuning MPC [22] uses closed-loop feedback to optimize controller parameters. However, it typically involves time-consuming data collections or requires actual events to trigger. In contrast, DT-MPC achieves runtime MPC tuning by inferring collisions from instantaneous Doppler measurements and obstacle state tracking, boosting real-time robustness for handling rapid-motion obstacles.
@@ III Problem Formulation
where $C_{t}(\mathcal{S}_{t})$ is the utility function, $d_{\text{safe}}$ is the safety distance, and $(\mathbb{G}_{h+1},{\mathbb{O}}_{h+1}^{n})$ are bounding boxes of ego-robot and the $n$-th obstacle, respectively. The minimum distance function between two bounding boxes is defined as $\mathbf{dist}(\mathbb{P},\mathbb{Q})=\min\{\|\mathbf{p}-\mathbf{q}\|_{2}\,|\,\mathbf{p}\in\mathbb{P},\mathbf{q}\in\mathbb{Q}\}$. Considering a reference path tracking task with target waypoints $\{\mathbf{s}_{h+1}^{\star}|h\in\mathcal{H}_{t}\}$, we formulate $C_{t}(\mathcal{S}_{t})=\sum_{h\in\mathcal{H}_{t}}\left\|\mathbf{s}_{h+1}-\mathbf{s}_{h+1}^{\star}\right\|_{2}^{2}$ [27].
Problem $\mathsf{P}_{t}$ is nontrivial since estimating ${\mathbb{O}}_{h+1}^{n}$ results in an inevitably biased $\hat{\mathbb{O}}_{h+1|t}^{n}$, which can significantly mismatch the actual ${\mathbb{O}}_{h+1}^{n}$ if obstacle $n$ has rapid motions. This error would further propagate to motion planning since $\mathbf{dist}(\mathbb{G}_{h+1},\hat{\mathbb{O}}_{h+1|t}^{n})\neq\mathbf{dist}(\mathbb{G}_{h+1},{\mathbb{O}}_{h+1}^{n})$.
Figure 3: Doppler velocity rectification.
@@ IV-A Doppler KalmanNet for Tracking
The ego robot uses a Doppler LiDAR to scan the surrounding environment and obtain Doppler point cloud $\mathcal{P}_{t}=\{\mathbf{c}_{t}^{i},\|\bm{\mu}_{t}^{i}\|_{2}\}_{i=1}^{I}$, where $I$ is the number of points in the frame, $\mathbf{c}^{i}_{t}\in\mathbb{R}^{3}$ and $\|\bm{\mu}^{i}_{t}\|_{2}\in\mathbb{R}$ are the spatial coordinate and the scalar Doppler velocity of the $i$-th point, respectively. Accordingly, we define the Doppler-augmented state $\mathbf{x}_{t}^{n}$ as
@@
where $(x_{t}^{n},y_{t}^{n})$, $\theta_{t}^{n}$, $\bm{v}_{t}^{n}$ and $\bm{a}_{t}^{n}$ represent the position, orientation, linear velocity and acceleration at the center, respectively. Each $\mathbf{x}_{t}^{n}$ can be used to generate a ${\mathbb{O}}_{t}^{n}$. Hence, we only need to estimate a sequence of $\{\hat{\mathbf{x}}_{t+1|t}^{n},\hat{\mathbf{x}}_{t+2|t}^{n},\ldots,\hat{\mathbf{x}}_{t+H|t}^{n}\}$ from $\mathcal{P}_{t}$. We can fuse $\mathcal{P}_{t}$ with a state transition model to achieve this goal, but two challenges exist. First, according to Doppler LiDAR manufacturers [15], real-world velocity measurements typically involve noises with standard deviation of $0.1$ m/s. Second, built upon past trajectories, the transition model may mismatch the actual obstacle motions in the future [17]. Our D-KalmanNet can simultaneously tackle both challenges.
IV-A1 Doppler Velocity Rectification
$\bm{\breve{\mu}}_{t}^{i}=\frac{\|\bm{\mu}_{t}^{i}\|_{2}}{\cos\rho_{t}^{i}}\times\frac{(\mathbf{I}-\mathbf{e}\mathbf{e}^{\top})(\mathbf{c}^{\text{D}}_{t}-\mathbf{c}_{t}^{i})}{\|(\mathbf{I}-\mathbf{e}\mathbf{e}^{\top})(\mathbf{c}^{\text{D}}_{t}-\mathbf{c}_{t}^{i})\|_{2}},$ (3)
@@ IV-A1 Doppler Velocity Rectification
4:   Group $\mathcal{P}_{t}^{n}$ to obtain $\mathbf{y}_{t}^{n}$
5:   if $t=0$ then   // Initial Guess
6:    $\mathbf{x}_{0}^{n}\leftarrow\mathbf{y}_{0}^{n}$, $\hat{\mathbf{x}}_{1|0}^{n}\leftarrow\mathbf{T}\cdot\mathbf{x}_{0}^{n}$, $\hat{\mathbf{y}}_{1|0}^{n}\leftarrow\mathbf{U}\cdot\hat{\mathbf{x}}_{1|0}^{n}$
7:   else   // Kalman Learning
8:    $\mathcal{K}_{t}^{n}\leftarrow\textsf{RNN}(\mathbf{x}^{n}_{t-1},\hat{\mathbf{x}}_{t|t-1}^{n},\hat{\mathbf{y}}_{t|t-1}^{n},\mathbf{y}_{t}^{n})$
9:    $\mathbf{x}_{t}^{n}\leftarrow\hat{\mathbf{x}}_{t|t-1}^{n}+\mathcal{K}_{t}^{n}\cdot(\mathbf{y}_{t}^{n}-\hat{\mathbf{y}}_{t|t-1}^{n})$
10:    $\hat{\mathbf{x}}_{h+1|t}^{n}\leftarrow\mathbf{T}^{h-t+1}\cdot\mathbf{x}_{t}^{n}$, $\forall h\in\mathcal{H}_{t}$
11:    $\hat{\mathbf{y}}_{t+1|t}^{n}\leftarrow\mathbf{U}\cdot\hat{\mathbf{x}}_{t+1|t}^{n}$
12:   end if
13:   return $\hat{\mathcal{O}}_{t}^{n}=\{\hat{\mathbb{O}}_{h+1|t}^{n}(\hat{\mathbf{x}}_{h+1|t}^{n})\}_{\forall h\in\mathcal{H}_{t}}$
@@
15:  $t\leftarrow t+1$
16:end while